OpenStudy (anonymous):

what it the equation of a line that has a slope of 2, a y intercept of 2 and passes through the points 1,1 and 2,0

OpenStudy (mathstudent55):

All you need is the slope and the y-intercept. Use the slope-intercept form of the equation of a line: \(y = mx + b\) where m = slope, and b = y-intercept.

OpenStudy (mathstudent55):

The line with slope 2 and y-intercept 2 does pass through the point (0, 2), but it does not pass through the point (1, 1). Are you sure you got that point correctly?

OpenStudy (anonymous):

sorry the slope is one

OpenStudy (anonymous):

a line which passes through (1,1) and (2,0) has slope \[=\frac{ 0-1 }{ 2-1 }=-1\] so check your question again.

OpenStudy (mathstudent55):

If the slope is 1 and the y-intercept is 2, then the equation is \(y = x + 2\) This line passes through (0, 2), but not through (1, 1).

OpenStudy (anonymous):

point is given (2,0) not (0,2)

OpenStudy (anonymous):

i think slope should be -1

OpenStudy (anonymous):

options y=-1/2x-2 y=-x-2 y=-1/2x+2 y=-x+2

OpenStudy (dan815):